find the distance between (2/3,4) and (5/3,7)
a- 10.06
b- 3.16
c- 3.46
d- 4.24

Answer by checkley75(3666)

THE X DIMENTION IS 5/3-2/3=3/3=1
THE Y DISTANCE IS 7-4=3
UXING THE PYTHAGOREAN THEROM WE GET
1^2+3^2=D^2
1+9=D^2
10=D^2
D=SQRT10
D=3.16 ANSWER (b)FOR THE DISTANCE BETWEEN THESE TWO POINTS.